Find the length of each segment if line segment AB and line segment CD are the same length.
Line Segment AB= 4(x+3)
Line Segment CD= 9x - 4

Answer by Alan3354(30993)

Line Segment AB= 4(x+3) = 4x + 12
Line Segment CD= 9x - 4
----------------
4x+12 = 9x-4
16 = 5x
x = 16/5
-----------
AB = 64/5 + 12 = 24.8
CD = 9x - 4 = 24.8